OpenStudy (anonymous):

What are the x-intercept(s) of the graph of y + 6 = x^2 − x? A.) (−2, 0) and (3, 0) B.) (−2, 0) and (−3, 0) C.) (2, 0) and (−3, 0) D.) (2, 0) and (3, 0)

OpenStudy (jack1):

rearrange the equation to the form y = ax^2 + bx +c first

OpenStudy (jack1):

x intercepts would occur at y=o so replace y with 0 therefore 0= ax^2 + bx +c

OpenStudy (anonymous):

Firstly you can subtract 6 from both sides of the equation leaving you with: y = x^2-x-6 (x ...)(x ...) = 0 This is a trick i learned once. The values on the ... must be so such that. the product of bots equals c (-6) and the sum of both equals b (-1) When we try a few values we find: (x-3)(x+2)=0 This equation only fits if one of these ( (x-3) or (x+2) ) equals zero. This is only the case if x = 3 or x = -2. And of course it only intercepts the x plane is y = zero. So the right answer here is (x,y) = (-2,0) and (3,0)

OpenStudy (anonymous):

oh my goodness thank you both!! ur the best!!!!